# Spica: Retrieving Scenarios for Pluralistic In-Context Alignment

### v1

@@ Abstract
Alignment of large language models (LLMs) to societal values should account for pluralistic values from diverse groups. One technique uses in-context learning for inference-time alignment, but only considers similarity when drawing few-shot examples, not accounting for cross-group differences in value prioritization. We propose Spica, a framework for pluralistic alignment that accounts for group-level differences during in-context example retrieval. Spica introduces three designs to facilitate pluralistic alignment: scenario banks, group-informed metrics, and in-context alignment prompts. From an evaluation of Spica on an alignment task collecting inputs from four demographic groups ($n=544$), our metrics retrieve in-context examples that more closely match observed preferences, with the best prompt configuration using multiple contrastive responses to demonstrate examples. In an end-to-end evaluation ($n=80$), we observe that Spica-aligned models are higher rated than a baseline similarity-only retrieval approach, with groups seeing up to a +0.16 point improvement on a 5 point scale. Additionally, gains from Spica were more uniform, with all groups benefiting from alignment rather than only some. Finally, we find that while a group-agnostic approach can effectively align to aggregated values, it is not most suited for aligning to divergent groups. [^1] We provide our code and data for others to build on: https://github.com/Social-Futures-Lab/SPICA-code
@@ 1 Introduction
The widespread availability of generative AI systems has increasingly highlighted situations where outputs can be inappropriate, misleading, or dangerous to users Weidinger et al. (2021); Ji et al. (2023); Qi et al. (2024). Correspondingly, researchers have explored embedding human values into generative models through various value alignment strategies Huang et al. (2024); Gabriel (2020); Christian (2021); Ouyang et al. (2022). Since these models are typically deployed to the general public, model creators and service providers often seek to align towards a one-size-fits-all set of universal values Bai et al. (2022). However, different groups within society often disagree on values and have different norms around when and how to apply values Gordon et al. (2022); Weld et al. (2022); Park et al. (2024). Recognizing this challenge, more recent work has approached model alignment from a pluralistic perspective Sorensen et al. (2024b); Feng et al. (2024)—rather than attempt to bridge irreconcilable differences, a better solution may be to support different perspectives of each group directly.
One general strategy for large language model (LLM) alignment—in-context alignment (ICA) Lin et al. (2024); Han (2023)—acts dynamically at inference time by retrieving few-shot examples of prompts and associated preferable responses as context. ICA is a promising strategy for pluralistic value alignment as it flexibly allows the encoding of different groups’ values through swapping in and out each group’s own preferred responses. However, simply extending existing ICA to pluralistic preference data is insufficient—in practice, different groups differ not only on values, but also norms that dictate which values are more salient Weld et al. (2022). As current ICA retrieval metrics only consider relevance of the prompt and not the group, the in-context demonstrations can end up only illustrating less important values despite showcasing similar prompts (Figure 1).
In this work, we present Spica, an evolution of retrieval-based in-context alignment, enabling alignment to a plurality of group-specific values while being cognizant of group-specific norms around such values. Spica consists of three main components: (1) scenario banks—shared collections of scenarios (prompts, responses, and group preferences) that can encode both values and norms; (2) group-informed retrieval measures—metrics that allow us to recover second-order norms from individual preference assessments; (3) ICL prompt setups that can effectively apply richer information from scenarios to the task of alignment.
We evaluated Spica by conducting an alignment task for four demographicgroups and examined three main aspects of the process: the quality of the scenarios retrieved, the effectiveness of different in-context prompts in applying scenarios to alignment, and the performance on the end-to-end task of alignment of model outputs.
@@
  • Compared to a baseline using only similarity-based scoring of examples, group-informed metrics retrieved scenarios that aligned more accurately (as measured by RMSE of ratings) to observed ground truth, indicating a quality gap caused by the reliance on similarity.
  • Among different ICL setups for integrating scenarios into prompts, the most effective designs were: P-I style—provide a single positive instruction when user preferences are collected over descriptions of response strategies; and C-R style—provide a contrasting spectrum of example responses when user preferences are collected over model outputs.
  • In an end-to-end evaluation, we find that Spica improves alignment overall more than baseline ICA (+0.053 / 5 points), with statistically significant gains (+0.16 / 5 points) observed on groups traditionally disadvantaged in alignment.
  • We also find that baseline ICA can result in disparate outcomes, with some groups preferring original non-aligned model outputs, whereas Spica alignment produces outputs uniformly preferred by all groups.
  • Finally, on alignment tasks for aggregated (collective) values, existing group-agnostic approaches can still perform well, despite not generalizing to pluralistic settings.
@@ 2 Related Work
Customizing LLMs for Value Alignment
Traditional methods for customizing LLMs for specific tasks and domains involve modifying training procedures. These include pretraining on task-specific corpora Wu et al. (2023); Lee et al. (2020), post-hoc finetuning Gururangan et al. (2020); Han and Eisenstein (2019), instruction tuning Ge et al. (2023); Gupta et al. (2022); Shi et al. (2023), and aligning with human preferences Ouyang et al. (2022). These approaches are also used to encode moral values and diverse human preferences into models Tay et al. (2020); Bai et al. (2022); Liu et al. (2022); Bang et al. (2023); Jang et al. (2023). However, they have significant limitations for value alignment. They require extensive human annotation to provide meaningful signals about desired values Kim et al. (2023), and even then, there is limited understanding or guarantee of how well the models have internalized these values Agarwal et al. (2024). This makes the models less robust in terms of value alignment. Moreover, once trained, these models lack flexibility; updating the model to reflect evolving values often requires a complete retraining, which is computationally intensive Carroll et al. (2024).
Retrieval-based In-Context Learning for Alignment
In-Context Learning (ICL) offers promising alternatives for value alignment by enabling behavior modifications during inference rather than training through the use of few-shot examples incorporated into model prompts Dong et al. (2022); Wei et al. (2022). The use of example demonstrations in ICL has also allowed systems to incorporate retrieval Lewis et al. (2020); Borgeaud et al. (2022) as a part of dynamically constructing in-context prompts informed by inputs Zhang et al. (2022); Rubin et al. (2021).
For the task of model alignment, approaches to using retrieval and in-context learning prompts, such as URIAL Lin et al. (2024), have also been referred to as in-context alignment (ICA) Han (2023). As most ICA systems today focus on addressing more collective preferences, the retrieval aspect of such systems has largely remained unchanged, with relatedness metrics like semantic similarity being the main way to rank retrieved examples Karpukhin et al. (2020); Gao et al. (2023). However, given examinations of how models apply in-context examples Min et al. (2022), there is room for improvement around how examples are ranked for retrieval, especially for value alignment tasks. Prior works around alignment have pointed to potentially important factors that could improve the utility of retrieved examples, such as noting potential importance for prioritizing examples that illustrate exceptional circumstances and edge cases Kiehne et al. (2022), or emphasizing examples that capturing population-specific preferences Hovy and Yang (2021); Kirk et al. (2023), with potentially more signals to consider for pluralistic alignment, where different groups can have different norms Weld et al. (2022) that moderate how preferences are prioritized over each other.
@@
Supporting pluralistic values is crucial for building general-purpose agents and LLMs Sorensen et al. (2024b) that are deployed into society. Large datasets like ValuePrism Sorensen et al. (2024a) and PRISM Kirk et al. (2024) highlight the importance of reflecting diverse values, yet achieving consensus remains challenging. Some approaches turn to higher-level abstract descriptions of values as a solution for improving consensus Bai et al. (2022), however, practical application to specific cases often reveals discrepancies Koshy et al. (2023). Drawing from the legal realm, there have also been approaches that propose combining higher-level descriptions with specific examples (e.g., legal precedents) to illustrate more ambiguous concepts encoded by values Cheong et al. (2024); Chen and Zhang (2023).
Beyond first-order challenges of encoding values themselves as mentioned above, pluralism can also give rise to second-order challenges when groups share similar sets of preferences or values (such as preferring diversity and factual quality) while also disagreeing on their salience Jackson (1960) (and thus prioritization) when it comes to practical application Weld et al. (2022). This aspect, however, is often overlooked by existing frameworks for pluralistic alignment.
@@ 3.1 Scenario Banks for Encoding Pluralistic Values and Norms
Past examples of data for alignment has included both normative data in the form of guidelines created through constitutional alignment processes Bai et al. (2022) and quantitative data in the form of user ratings of conversations between humans and LLMs Kirk et al. (2024). While this data can be incorporated in ICA settings via prompts and example demonstrations respectively, it can be insufficient for the needs of pluralistic alignment. It can be difficult to attribute normative guidelines created through deliberation and consensus to each possible group in a population. Additionally, human ratings of conversations often only apply to existing outputs making it difficult to evaluate the broader space of responses.
@@
Given this, with Spica, we propose a way of collecting pluralistic alignment data in the form of scenario banks. A scenario consists of three main components: (1) a prompt ($x$), which represents an example of a user query or conversation with a model leading up to a morel response; (2) responses ($y\in Y_{x}$), the space of possible ways a model could respond to a prompt, which can take the form of either specific examples of outputs, or high-level strategies covering a class of outputs; and (3) preferences ($r_{p}(x,y)$), which encode an individual $p$’s preference of a response $y$ to a prompt $x$. A scenario bank consists of a collection of such scenarios and provides a basis for the ground truth in an ICA retrieval. By creating scenario banks this way, we are able to understand distributional properties of preferences both at an individual level and at a group level, allowing us to understand both individual values and group norms in pluralistic alignment.
First, encoding values as distributions over responses provides better understanding to human preferences in context. Traditional human ratings collected for disjoint prompt-response pairs lead to rating values that need to be compared between prompts—e.g., there is not sufficient information to tell if a particular low-rated response to a prompt was just a bad response, or actually the “best possible” out of a prompt with no good options. Even when multiple responses from different models are evaluated against each other Kirk et al. (2024), we are still limited by distributional properties of the pre-alignment outputs which could lead to groups less represented by the training data struggling to indicate desirable outputs to them.
Second, by focusing on individuals, we can recover norms by looking at distributional properties within groups—even when groups may be (re-)constructed post-hoc. Theory around the structural nature of social norms Jackson (1960) pose that norms can be fundamentally observed as patterns in preferences that are salient across individuals within a group, which can now also be measured through scenario banks. Additionally, while other normative approaches like directly asking for high-level criteria can also uncover norms, outputs can be hard to disentangle as individuals can often have intersectional identities that span multiple groups. With scenario banks, one only needs to find the individuals corresponding to a group should group criteria change or new groups be considered, adding new data only if insufficient.
Finally, distributional data allows us better understanding of alignment outcomes, which can be especially important due to the stochastic nature of outputs. Given a ground truth rating distribution (assuming discrete response classes) $r(x,y),\forall y\in Y_{x}$, we can compare different alignments by simply observing the amount of divergence between the two. In this discrete case, we can simply use a loss-style function like the root mean squared error (RMSE):
@@
We can use a similar approach to understand retrieval in ICA, where we can find the expected loss across a group if the model produced outputs similar to the retrieved $x^{\prime}$ but for $x$: $\mathbb{E}_{r}[L(r(x^{\prime}),r(x))]$.
@@ 3.2 Group-Informed Retrieval Measures
As noted earlier, in pluralistic alignment, values alone are often insufficient as communities Weld et al. (2022) and demographic groups Kumar et al. (2021) can have similar preferences on specific examples while making different higher-level trade-offs around what are salient examples—e.g., across many examples one may find that a group is prioritizing correctness over respectfulness, or helpfulness over safety, even when each property assessed individually comes out as desirable. Existing retrieval metrics which only compare similarity of the input prompts $x$ and known examples $x^{\prime}$ have no way to encode this group-level difference.
For a solution to this, we take inspiration from prior work that examines structural properties of social norms, with one such example being the return potential model Jackson (1960). This model interprets norms as distributional properties over behavior appropriateness that arise when viewing a group in aggregate. In the alignment setting, we adapt two concepts to assess model outputs (behavior): (1) We adapt the idea of “crystallization”—whether preferences over a behavior is consistent across group members—into the metric $g_{\text{stability}}(x^{\prime})$; and (2) we adapt the ideas of “intensity” and “tolerable range”—which assess differences in preference across behaviors for the same group member—into the metric $g_{\text{contrast}}(x^{\prime})$.
3.2.1 Stability: Differentiating Common Values from Individual Preferences
With social norms, “crystallization” assesses whether preferences on any particular behavior has become shared among the group, and thus become more crystallized as a norm. In the ICA setting, this concept can also be useful: While all ICL examples have the potential to illustrate preferences within a group, an example that illustrates a set of shared (crystallized) preferences is likely to be a better example demonstration for the group. We measure this for each scenario in our scenario bank: if, for that scenario $x^{\prime}$, the variance between annotators’ preferences $r_{p}(x^{\prime},y^{\prime})$ on each response type $y^{\prime}\in Y_{x^{\prime}}$ is lower, then the scenario is likely to demonstrate more crystallized norms than weaker preferences. More concretely:
@@ 3.2.2 Contrast: Assessing Indifference versus Preference
With social norms, concepts like “tolerable range” and“intensity” assess how broad the range of acceptable (and unacceptable) behaviors is and the intensity at which individuals express this preference Hackman (1992). Like with crystallization, we can also assess a similar concept with scenario banks: For a scenario $x^{\prime}$, the variance between different behaviors $y^{\prime}\in Y_{x^{\prime}}$ across each annotator $r_{p}(x^{\prime},y^{\prime})$ assesses how much they care about differentiating preferences. More concretely:

### v2

@@ Abstract
When different groups’ values differ, one approach to model alignment is to steer models at inference time towards each group’s preferences. However, techniques like in-context learning only consider similarity when drawing few-shot examples and not cross-group differences in values. We propose Spica, a framework that accounts for group-level differences during in-context example retrieval. Spica introduces three designs: scenario banks, group-informed retrieval metrics, and in-context alignment prompts. From an evaluation of Spica on an alignment task collecting inputs from four demographic groups ($n=544$), our metrics retrieve in-context examples that more closely match observed preferences, with the best prompt configuration using multiple contrastive responses to demonstrate examples. In an end-to-end evaluation ($n=120$), we observe that Spica is higher rated than similarity-based retrieval, with groups seeing up to a +0.16 point improvement on a 5 point scale. Additionally, gains from Spica were more uniform, with all groups benefiting from alignment rather than only some. Finally, we find that while a group-agnostic approach can align to aggregated values, it is not most suited for divergent groups. [^1] We provide our code and data for others to build on: https://github.com/Social-Futures-Lab/SPICA-code
@@ 1 Introduction
The widespread availability of generative AI systems has highlighted how outputs can be inappropriate or dangerous to users Weidinger et al. (2021); Ji et al. (2023); Qi et al. (2024). Correspondingly, researchers have explored embedding human values into models through various alignment strategies Huang et al. (2024); Gabriel (2020); Christian (2021); Ouyang et al. (2022). Typically, model providers seek to align towards a one-size-fits-all set of universal values Bai et al. (2022). However, different groups within society often disagree on values and have different norms around when and how to apply values Gordon et al. (2022); Weld et al. (2022); Park et al. (2024). More recent work has called for a pluralistic perspective Sorensen et al. (2024b); Feng et al. (2024)—rather than try to bridge irreconcilable differences, we should directly support different perspectives of each group.
One general strategy for large language model (LLM) alignment—in-context alignment (ICA) Lin et al. (2024); Han (2023)—acts dynamically at inference time by retrieving few-shot examples of prompts and associated preferable responses as context. ICA is a promising strategy for steerable pluralistic alignment as different groups can use their own examples to illustrate their values. However, pluralistic alignment extends beyond illustrating different values—prior work has observed that across online communities, not only can collective values differ, norms around how important values are in relation to each other can also differ Weld et al. (2022). When considering ICA for pluralistic alignment, simply focusing on whether examples illustrate some relevant values is insufficient. It is also important to consider whether these examples demonstrate the salient ones given group or community norms (Figure 1).
In this work, we present Spica, an evolution of retrieval-based in-context alignment that focuses on pluralistically aligning model outputs to values and norms of different groups. Spica consists of three main components: (1) scenario banks—shared collections of scenarios (prompts, responses, and group preferences) that can encode both values and norms; (2) group-informed retrieval measures—metrics that allow us to recover second-order norms from individual preference assessments; (3) ICL prompt setups that can effectively apply richer information from scenarios to the task of alignment.
We evaluated Spica by conducting an alignment task where we take a base model and produce pluralistically aligned outputs for four demographicgroups . We examined three aspects of the process: the quality of the scenarios retrieved, the effectiveness of different in-context prompts in applying scenarios to alignment, and performance on the end-to-end task of alignment of model outputs.
@@
  • Compared to a baseline using only similarity-based scoring, group-informed metrics retrieved scenarios that aligned more accurately to observed ground truth, indicating a quality gap when only relying on similarity.
  • Among different prompting setups for integrating retrieved scenarios, the most effective designs were: P-I style—provide a single positive instruction when user preferences are collected over descriptions of response strategies; and C-R style—provide a contrasting spectrum of example responses when user preferences are collected over model outputs.
  • In an end-to-end evaluation, we find that Spica produces more aligned outputs than baseline ICA (+0.053 / 5 points), with statistically significant gains (+0.16 / 5 points) observed on traditionally disadvantaged groups.
  • We also find that baseline ICA can result in disparate outcomes, whereas Spica alignment produces outputs uniformly preferred by all.
  • Finally, we examine Spica’s group-informed metric on collective alignment settings, noting that for aggregate values, group-agnostic approaches tend to be sufficient.
@@ 2 Related Work
Value Alignment of LLMs
Traditional methods for customizing LLMs for specific tasks and domains involve modifying training procedures. These include pretraining on task-specific corpora Wu et al. (2023); Lee et al. (2020), post-hoc finetuning Gururangan et al. (2020); Han and Eisenstein (2019), instruction tuning Ge et al. (2023); Gupta et al. (2022); Shi et al. (2023), and aligning with human preferences Ouyang et al. (2022). These approaches are also used to encode moral values and human preferences Tay et al. (2020); Bai et al. (2022); Liu et al. (2022); Bang et al. (2023); Jang et al. (2023). However, they have significant limitations for value alignment. They require extensive human annotation to provide meaningful signals about desired values Kim et al. (2023), and even then, there is limited understanding of how well the models have internalized these values Agarwal et al. (2024), making them less robust for value alignment. Moreover, once trained, these models lack flexibility; updating the model to reflect evolving values often requires complete retraining Carroll et al. (2024).
In-Context Learning for Alignment
In-Context Learning (ICL) offers promising alternatives by enabling behavior modifications during inference rather than training through the use of few-shot examples incorporated into model prompts Dong et al. (2022); Wei et al. (2022). The use of example demonstrations in ICL has also allowed systems to incorporate retrieval Lewis et al. (2020); Borgeaud et al. (2022) as a part of dynamically constructing in-context prompts informed by inputs Zhang et al. (2022); Rubin et al. (2021).
For the task of model alignment, approaches to using retrieval and in-context learning prompts, such as URIAL Lin et al. (2024), have also been referred to as in-context alignment (ICA) Han (2023). As most ICA systems focus on addressing collective preferences, how they do retrieval has largely remained unchanged, with relatedness metrics like semantic similarity being the main way to rank retrieved examples Karpukhin et al. (2020); Gao et al. (2023). Prior works around alignment have suggested ways to potentially improve the utility of retrieved examples, such as prioritizing examples that illustrate exceptional circumstances and edge cases Kiehne et al. (2022), or emphasizing examples that capture population-specific preferences Hovy and Yang (2021); Kirk et al. (2023). These signals are further complicated in pluralistic settings, where different groups can have different norms Weld et al. (2022) that moderate how preferences are prioritized over each other.
@@
Supporting pluralistic values is crucial for building general-purpose agents and LLMs Sorensen et al. (2024b). Large datasets like ValuePrism Sorensen et al. (2024a) and PRISM Kirk et al. (2024) highlight the importance of reflecting diverse values, yet achieving consensus remains challenging. Some approaches turn to higher-level abstract descriptions of values as a solution for building consensus via deliberative inputs Bai et al. (2022). However, practical application of these values to specific cases often reveals discrepancies in understanding Koshy et al. (2023). Drawing from the legal realm, there have also been approaches that propose combining higher-level descriptions with specific examples (e.g., legal precedents) to illustrate more ambiguous concepts encoded by values Cheong et al. (2024); Chen and Zhang (2023).
Beyond first-order challenges of encoding values, pluralism can also give rise to second-order challenges when groups share similar sets of preferences or values (such as preferring diversity and factual quality) while also disagreeing on their salience Jackson (1960) and thus prioritization in practical application Weld et al. (2022). This aspect is often overlooked by existing frameworks for pluralistic alignment. Spica addresses this by capturing disaggregated individual preferences that can be used to derive both first-order group preferences (values) and second-order group norms.
@@ 3.1 Scenario Banks for Encoding Pluralistic Values and Norms
Past examples of data for alignment have included both normative data in the form of “constitutional” guidelines Bai et al. (2022) and quantitative data in the form of user ratings of conversations between humans and LLMs Kirk et al. (2024). While both types of input can be used for ICA, pluralistic alignment introduces additional challenges. Normative guidelines require deliberation to create, which can be costly if each group needs to (re-)convene each time to make their own. On the other hand, ratings of conversations are limited by the behavior of existing models, making it challenging for groups to significantly deviate on norms or values.
@@
Given this, with Spica, we propose a way of collecting pluralistic alignment data in the form of scenario banks, which uses prompts and responses guided by classes of model behaviors to ground the collection of dis-aggregated ratings, addressing the limitations above. A scenario consists of three main components: (1) a prompt ($x$)—an example of a user query or conversation with a model leading up to a response; (2) responses ($y\in Y_{x}$)—the space of possible ways a model could respond to a prompt, which can take the form of either specific examples of outputs, or high-level response classes covering many outputs; and (3) preferences ($r_{p}(x,y)$)—ratings that encode an individual $p$’s preference of a response $y$ to a prompt $x$. A scenario bank consists of a collection of such scenarios and provides a basis for the ground truth in an ICA retrieval. Through dis-aggregated data and classes of behaviors, scenario banks allow us to recover group values by taking consensus across individuals, and understand group norms by observing distributions of ratings across individuals.
With scenarios, individual preferences are collected as distributions of ratings over a known span of response classes. This provides more contextual understanding of preferences—e.g., did a user rate a response lowly because it was a less appropriate way to respond, or are other ways of responding even worse? Existing evaluations of responses from different models to the same query Kirk et al. (2024) do provide similar distribution-level information. However, existing model outputs have been shown to be biased Buyl et al. (2024); Rozado (2024), which can make it hard for groups to indicate their values this way due to lack of outputs that follow their values.
Additionally, the dis-aggregated nature of individual ratings means that we are not limited to the consensus of normative guidelines, and can instead reconstruct values and norms post-hoc. Shared values can be constructed by taking consensus preferences from individual ratings, while group-level social norms can be observed from how individuals within a group agree or disagree with one another Jackson (1960).
3.1.1 Comparing Preferences over Model Behaviors
In scenario banks, preferences are collected as rating distributions across a set of model behaviors, reflected as classes of responses (see Table 1). Using this formulation, we can compare not only the preferences for any specific behavior, but also how well preference distributions (of users or models) align with each other. Taking any two preference distributions $r(x,y)$ and $r^{\prime}(x,y)$, we can define how much they diverge by observing how much they disagree across the different response classes $y\in Y_{x}$, which we can measure with a loss based on the root mean squared error (RMSE):
@@ 3.1.1 Comparing Preferences over Model Behaviors
Here, $r$ could reflect an individual’s preference, a consensus preference from a group, or even a retrieval-based ICA model’s implied “preference”—by retrieving $x^{\prime}$ as a demonstration for $x$, the model implies that it expects $r(x,y)$ to match $r(x^{\prime},y^{\prime})$ for corresponding response classes $y$ and $y^{\prime}$.
@@ 3.2 Group-Informed Retrieval Measures
As noted earlier, values alone are often insufficient as communities Weld et al. (2022) and demographic groups Kumar et al. (2021) can have similar values (seen as preferences on specific examples) while making different higher-level trade-offs around what are salient examples when—e.g., across many scenarios, one may find that a group is prioritizing correctness over respectfulness, or helpfulness over safety, even when they view all these properties as positive in isolation. Existing retrieval metrics only compare similarity of the input prompts $x$ and known examples $x^{\prime}$ and do not encode this group-level difference. To address this, we take inspiration from the return potential for social norms Jackson (1960), and define two group-informed measures. First, we adapt the idea of “crystallization”—whether particular values are consistently held across group members—into the metric $g_{\text{stability}}(x^{\prime})$. Second, we adapt the ideas of “intensity” and “tolerable range”—whether individuals in the group are more opinionated or ambivalent—into the metric $g_{\text{contrast}}(x^{\prime})$. Together, these metrics interpret the distributional nature of individual values (preferences over model behavior) within a group to identify emergent norms.
3.2.1 Stability: Differentiating Norms from Individual Values
With social norms, “crystallization” describes whether a behavior preference (value) is consistently held across different members in the group such that it has become crystallized as a norm. We borrow this concept for ICA to assess group norms: For some example scenario, by looking at preferences across members within the group on each model behavior, we can assess whether members tend to agree, which would indicate the scenario reflects a norm, or disagree, which indicates a less salient example. More formally: if, for a potentially retrieved scenario $x^{\prime}$, the variance between annotators’ preferences $r_{p}(x^{\prime},y^{\prime})$ on each response type $y^{\prime}\in Y_{x^{\prime}}$ is lower, then the scenario is likely to demonstrate more crystallized norms than weaker preferences.
@@ 3.2.2 Contrast: Assessing Indifference versus Preference
With social norms, concepts like “tolerable range” and “intensity” assess how broad the range of acceptable (and unacceptable) behaviors is and the intensity at which individuals express this preference Hackman (1992). In the context of ICA, examples that illustrate stronger preferences for sets of behaviors are more valuable than those that simply indicate indifference. Here we can also create a metric based on the dis-aggregated preferences from scenario banks: For a scenario $x^{\prime}$, the variance between different behaviors $y^{\prime}\in Y_{x^{\prime}}$ across each annotator $r_{p}(x^{\prime},y^{\prime})$ assesses how much they care about differentiating preferences. More concretely: